A sum of $10000 is split, and the two parts are invested at 5% and 6% respectively. if the interest from the 5% investment exceeds the interest from the 6% investment by $60, how much is invested at each rate?
Found 2 solutions by Nate, pigpen:
Answer by Nate(3500)

The answer depends on the amount of time....

Answer by pigpen(1)

.05(10000-x)=.06x+60
500-.05x = .06x+60
+.05x +.05x
500 = .11x+60
-60 -60
440 = .11x
--- ----
.11 .11
4000 = x
$4000 @ 5%
10000 - x = 6% so 10000-4000 = $6000 @6%
pigpen